In Problems 9 to 12, graph the function that the requested Trigonometric Fourier Series of the indicated problem above conver5. f(x)= In Problems 5 to 8, determine the Fourier Cosine Series that converges to the given function at each of its points oHomework Problems for Handout Sheet 25 In Problems 1 to 4, determine the Fourier Sine Series that converges to the given funcPlease help in any of number 9-12 in differential equations

the questions 9-12 uses the information from the other pictures posted of questions 1-8
Answer #1

Solution for 10. Following the solution to 3 to obtain fourier sine series. Then I have plotted using desmos for 100 terms. Graph is for - 4pi to 4pi. 4pi=12.56637

3. Fourier sine series f(x) = 25-x on > This means period = Wit i flatuit)=f(x) 0<xsat zit D 2t 2 - 12 JJ B = 255(x) sin norxΓ -10 Σε sin(1) η = 1 4π = 12.5663706144 powered by desmos